# LeMiCa: Lexicographic Minimax Path Caching for Efficient Diffusion-Based Video Generation

### v1

@@ B.1 Models
In this paper, we introduce LeMiCa, a novel caching technique designed to accelerate and enhance a range of state-of-the-art video synthesis models, including Open-Sora 1.2 (Open-Sora, ), Latte (ma2025latte, ), and CogVideoX (yang2024cogvideox, ). Open-Sora 1.2 integrates 2D/3D VAEs and ST-DiT blocks for efficient video compression and generation. Latte leverages spatio-temporal tokenization and Transformer layers to model video distributions in the latent space. CogVideoX employs a 3D VAE and expert Transformers with adaptive LayerNorm for modality fusion and high-fidelity generation. In our experiments, we adopt the CogVideoX-2B variant.
@@ Appendix F Extension to text-to-image model
To further evaluate the scalability of our approach, we extend LeMiCa to the commercial text-to-image model Qwen-Imagewu2025qwenimagetechnicalreport . As shown in Figure 13 and Table 8, LeMiCa achieves strong generalization while maintaining high efficiency and visual fidelity, consistently outperforming TeaCache and DBCachecache-dit@2025 across both qualitative and quantitative metrics.
@@
Table 8: Comparison of inference efficiency and visual quality of LeMiCa and other acceleration methods on the Qwen-Image.
@@
Original: 50 steps 32.68 1.00× 1.18 33.63
30% steps 9.86 3.31× 1.10 31.62 0.36 0.73 15.83
TeaCache (l=0.6) 18.52 1.76× 1.09 32.60 0.42 0.70 14.90
DBCache (Fn8, Bn0, r0.6) 13.95 2.34× 0.87 32.62 0.36 0.78 21.45
LeMiCa (B=17) 13.31 2.46× 1.16 33.39 0.09 0.93 27.92
DBCache (Fn4, Bn2, r1.5) 9.57 3.41× -2.06 15.50 0.89 0.13 5.56
DBCache + TS (O=4)* 9.70 3.37× -0.23 29.75 0.63 0.65 16.57
LeMiCa (B=10) 9.75 3.35× 1.15 33.49 0.27 0.80 18.58
* DBCache is derived from the Cache-DiT acceleration framework, and TS denotes the Taylorseer acceleration method.

### v2

@@ 5 Conclusion
LeMiCa: Lexicographic Minimax Path Caching for Efficient Diffusion-Based Video Generation
@@ B.1 Models
In this paper, we introduce LeMiCa, a novel caching technique designed to accelerate and enhance a range of state-of-the-art video synthesis models, including Open-Sora 1.2 (Open-Sora), Latte (ma2025latte), and CogVideoX (yang2024cogvideox). Open-Sora 1.2 integrates 2D/3D VAEs and ST-DiT blocks for efficient video compression and generation. Latte leverages spatio-temporal tokenization and Transformer layers to model video distributions in the latent space. CogVideoX employs a 3D VAE and expert Transformers with adaptive LayerNorm for modality fusion and high-fidelity generation. In our experiments, we adopt the CogVideoX-2B variant.
@@ Appendix F Extension to text-to-image model
To further evaluate the scalability of our approach, we extend LeMiCa to the commercial text-to-image model Qwen-Imagewu2025qwenimagetechnicalreport. As shown in Figure 13 and Table 8, LeMiCa achieves generalization while maintaining high efficiency and visual fidelity, with competitive performance in visual perception metrics compared to Cache-DiTcache-dit@2025.
@@
Table 8: Comparison of inference efficiency and visual quality of LeMiCa and Cache-DiT on the Qwen-Image.
@@
Original: 50 steps 32.68 - 1.196 35.116
Cache-DiT (F4W4R0.32) 13.48 2.42x 1.174 35.067 0.137 0.885 32.254
Cache-DiT (F1W8R0.8) 11.24 2.91x 1.166 35.245 0.193 0.853 30.373
Cache-DiT (F1W4R0.8) 9.74 3.35x 1.106 35.206 0.337 0.760 29.143
LeMiCa (B=17) 13.31 2.45x 1.163 35.132 0.100 0.929 34.326
LeMiCa (B=10) 9.69 3.37x 1.134 35.253 0.253 0.804 30.316